### v2

@@ 4.1 Prompt Development
Transform legal text in natural language to expressions
in Defeasible Deontic Logic (DDL) in XML format. Each
atom should end with "(X)". If you want to represent a
conjunction, separate the atoms by a comma. If you want
to represent a disjunction, please use multiple rules
and do not write it as a single rule. Output only a
single <Paragraph> element with multiple <Rule> elements
if necessary. Make sure to output valid XML. Represent
obligations with [O], permissions with [P], and
prohibitions with [F]. If you want to negate an atom,
use the negation symbol "-" before the atom or the
deontic operator. Each rule should have only one
consequence. If you want to represent multiple
consequences, please use multiple rules. Since the law
snippet are talking about complaint handling, in most
preconditions, there will be an atom like complaint(X).
Make sure to keep the atoms in the precondition as
simple as possible. If it is possible to break down the
atoms into smaller parts, please do so. For example,
instead of urgentComplaint(X), write complaint(X),
urgent(X). Moreover, NEVER put an atom in the antecedent
if it also appears in the consequence, because this
would be syntactically invalid.
@@
4. Formalize the rules in the given format using
Defeasible Deontic Logic (DDL).
@@ 4. Formalize the rules in the given format using
8.1.1 A Supplier must take the following actions to
enable this outcome:
(c) Ensure awareness and visibility: ensure their
staff who have direct contact with Consumers or
former Customers, including personnel working for
contractors, understand the Suppliers Complaint
handling process, their responsibilities under it
and are able to identify and record a Complaint.
@@ 8.1.1 A Supplier must take the following actions to
8.1.1 A Supplier must take the following actions to
enable this outcome:
(a) Implement a process: implement, operate and comply
with a Complaint handling process that:
@@
D. requiring Consumers or former Customers to be
advised of the Resolution of their Complaint; and
@@
complaint(X), resolution(X) =>
[O] informResolution(X)
@@
8.5.1 A Supplier must take the following actions to
enable this outcome:
(e) Maintain confidentiality: Suppliers not subject to
the requirements of the Privacy Act must ensure
personal information concerning a Complaint is not
disclosed except as required to manage a Complaint
with the TIO or with the express consent of the
Consumer.
@@ 8.5.1 A Supplier must take the following actions to
complaintHandlingProcess(X),
personalInformation(X),
-subjectPrivacyAct(X) =>
@@
personalInformation(X), requestFromTIO(X) =>
[O] discloseInformation(X)
@@
The final prompt was evaluated across multiple LLMs. Two diagrams summarize the results: one displaying the standard success scores $S$ (s. Figure 1) and another illustrating the success scores $S^{*}$ under the stricter criterion of perfect formalizations (s. Figure 2).
@@ 4.2 Consideration of Multiple Law Snippets Simultaneously
In the experiments described in Section 4.1, the prompt was sent together with an individual law snippet to the LLM. This approach minimized token consumption but limited the reuse of atom names across snippets. Here, we investigate whether providing multiple law snippets simultaneously enhances formalization performance.
@@ 4.2.1 Incorporating the Formalization History
However, no improvement was observed compared to the single-snippet baseline; in fact, the success scores were marginally lower. A plausible explanation is that the additional context overwhelmed the models, hindering their ability to focus effectively on the current snippet.
@@ 4.2.3 Formalizing All Law Snippets in a Single Interaction
Although this approach led to a slight increase in atom reuse across snippets, it exhibited a major drawback: the generated formalizations were often less detailed compared to the baseline obtained in Section 4.1. In particular, important facts were frequently merged into a single rule, even in cases where separate rules would have been necessary for a proper and precise formalization.
@@ 4.3 Fine-Tuning
Fine-tuning is a transfer learning technique where pretrained model weights are adapted to a new task through further training. By leveraging knowledge acquired during pretraining, fine-tuning can substantially enhance model performance, particularly in scenarios characterized by limited training data. Prior work has demonstrated the effectiveness of fine-tuning LLMs in improving task-specific outcomes [25].
In the present study, fine-tuning experiments were conducted with GPT-4o. Although the proprietary nature of GPT-4o does not allow direct access to the model weights, OpenAI provides fine-tuning capabilities for non-reasoning models via its platform. Note that OpenAI’s reasoning models are not amendable to fine-tuning.
Given that only 22 law snippets from the dataset presented in [8] correspond to Sections 8.2.1(a)–8.2.1(c), the remaining 44 snippets from unrelated sections were utilized as training data.
@@
Configuration 1 parameters were determined automatically by OpenAI, as recommended for initial fine-tuning attempts [26]. However, early signs of overfitting motivated adjustments such as increasing the batch size and reducing the learning rate in Configurations 2 and 3.
The resulting performance is depicted in Figures 5 and 6, where blue bars correspond to non-fine-tuned baselines and other colors represent fine-tuned models. After each training epoch, an evaluation has been conducted.
Fine-tuning resulted in an improved success score after a single epoch of training under Configurations 2 and 3, relative to the baseline performance of the non-fine-tuned GPT-4o. However, subsequent training epochs led to a decline in performance, indicative of overfitting.
@@
4.4 Two-Stage Pipeline
In this approach, a two-stage prompting strategy was employed, where the output of the first stage served as part of the input for the second stage. This method aligns with the Layer-Of-Thoughts paradigm, which has been shown to enable complex reasoning in LLMs [27].
In the first stage, atom names were extracted from the legal texts. To this end, the LLMs were instructed to identify atom names alongside brief textual descriptions (Listing 3). Three illustrative examples were provided within the prompt, thus applying a three-shot learning strategy.
Extract all the relevant atoms from the legal text in
natural language and add a textual description of them.
Each atom should end with "(X)". Do not include
negations in the atom name - these will be introduced
later on. Since the law snippet are talking about
complaint handling, in most law snippets, there will be
an atom like complaint(X). Make sure to keep the atoms
as simple as possible. If it is possible to break down
the atoms into smaller parts, please do so. For example,
instead of urgentComplaint(X), write complaint(X),
urgent(X). The only exception to this rule is when you
can anticipate that an atom will belong into the
consequence. In this case, a longer atom name is better,
as each rule can have only one consequence. Keep in mind
that these atoms will serve as antecedents and
consequences in formalized rules - therefore, formalize
enough atoms so that antecedents and consequents can be
constructed from them. Formalize at least two atoms.
@@ 4.4 Two-Stage Pipeline
8.1.1 A Supplier must take the following actions to
enable this outcome:
(a) Implement a process: implement, operate and
comply with a Complaint handling process that:
(v) clearly states that Consumers or former
Customers have a right to make a Complaint and
that a proposed Resolution must be accepted by a
Consumer or former Customer before a Supplier is
required to implement it;
@@ 8.1.1 A Supplier must take the following actions to
informRightToMakeComplaint(X): Supplier informs customer
of right to make a complaint.
informComplaintHandlingProcess(X): Supplier informs
Customer of Complaint handling process.
complaintHandlingProcess(X): Supplier has a complaint
handling process as per TCPC section 8.
@@
Four separate experiments on atom extraction were conducted, involving the models DeepSeek-R1, a fine-tuned variant of GPT-4o, OpenAI o3, and OpenAI o4-mini. As in Section 4.3, the 44 law snippets from [8] not associated with Sections 8.2.1(a)–8.2.1(c) were used as training data for the fine-tuning variant.
In the subsequent stage, the generation of DDL rules was performed based on the legal text and the previously extracted atom definitions.
@@
Figure 7: Success scores of the two-stage pipeline
@@
Figure 8: Success scores of the two-stage pipeline (perfect formalizations only)
The results indicate that employing DeepSeek-R1 for atom extraction yields the most favorable outcomes, although they are inferior to the best results obtained without using a two-stage pipeline.
Furthermore, since overall performance remained relatively stable regardless of the model used in the second stage, it can be inferred that the primary source of error originates from the atom extraction step. Consequently, further optimization efforts should prioritize improving atom extraction rather than refining the second stage of the pipeline.
4.5 Comparison with Dragoni et al. [8]
@@
In this section, we compare our experimental findings with the results presented by Dragoni et al. [8]. To ensure methodological consistency, we restrict the comparison to the lower branch evaluation reported in their study. For this purpose, we employ the standard metrics of Precision, Recall, and their harmonic mean, the F1-score, defined as follows:
@@
Our analysis reveals an immediate discrepancy: Dragoni et al. report 65 terms and 36 rules in the gold standard, whereas we identified 69 terms and 52 rules across Sections 8.2.1(a)–8.2.1(c) in the gold standard.
Importantly, in the following analysis, we evaluate the precision and recall of the formalizations produced by the LLMs based on our counts. As a result, the reported metrics are not directly comparable to those presented in the work of Dragoni et al.
@@ 4.5.1 Evaluation of Term Identification
The first level of comparison involves the correct identification of legal terms (referred to as atoms). Dragoni et al. report a precision of $83.05\%$ and a recall of $90.78\%$. Table 2 summarizes the precision and recall achieved across various configurations and models in our study. The best precision, $86.21\%$, is achieved with DeepSeek-R1 when all law snippets were provided together within a single user prompt. The best recall, $84.06\%$, is achieved in the baseline setting with GPT-o3.
Table 2: Precision and Recall for Term Identification
Experiment Model
Generated
Atoms
Matched
Atoms
Precision Recall F1
Results from [8] $83.05\%$ $\mathbf{90.78\%}$ $\mathbf{0.87}$
Simple CoI (Section 4.1) o3 $93$ $58$ $62.37\%$ $84.06\%$ $0.72$
o4-mini $96$ $57$ $59.38\%$ $82.61\%$ $0.69$
DeepSeek-R1 $84$ $56$ $66.67\%$ $81.16\%$ $0.73$
GPT-4o $77$ $56$ $72.73\%$ $81.16\%$ $0.77$
Single Interaction (Section 4.2.3) o4-mini $64$ $52$ $81.25\%$ $75.36\%$ $0.78$
DeepSeek-R1 $58$ $50$ $\mathbf{86.21\%}$ $72.46\%$ $0.79$
GPT-4o $64$ $54$ $84.38\%$ $78.26\%$ $0.81$
Fine-tuned (Section 4.3) GPT-4o (Config 1, Epoch 1) $57$ $42$ $73.68\%$ $60.87\%$ $0.67$
GPT-4o (Config 2, Epoch 1) $71$ $57$ $80.28\%$ $82.61\%$ $0.81$
GPT-4o (Config 3, Epoch 1) $72$ $57$ $79.17\%$ $82.61\%$ $0.81$
Two-Stage (Section 4.4) DeepSeek-R1 $\rightarrow$ DeepSeek-R1 $73$ $49$ $67.12\%$ $71.01\%$ $0.69$
DeepSeek-R1 $\rightarrow$ GPT-4o $72$ $49$ $68.06\%$ $71.01\%$ $0.70$

### v3

@@ 4.1 Prompt Development
Transform legal text in natural language to expressions in Defeasible Deontic Logic
(DDL) in XML format. Each atom should end with "(X)". If you want to represent a
conjunction, separate the atoms by a comma. If you want to represent a disjunction,
please use multiple rules and do not write it as a single rule. Output only a
single <Paragraph> element with multiple <Rule> elements if necessary. Make sure to
output valid XML. Represent obligations with [O], permissions with [P], and
prohibitions with [F]. If you want to negate an atom, use the negation symbol "-"
before the atom or the deontic operator. Each rule should have only one
consequence. If you want to represent multiple consequences, please use multiple
rules. Since the law snippet are talking about complaint handling, in most
preconditions, there will be an atom like complaint(X). Make sure to keep the atoms
in the precondition as simple as possible. If it is possible to break down the
atoms into smaller parts, please do so. For example, instead of urgentComplaint(X),
write complaint(X), urgent(X). Moreover, NEVER put an atom in the antecedent if it
also appears in the consequence, because this would be syntactically invalid.
@@
4. Formalize the rules in the given format using Defeasible Deontic Logic (DDL).
@@
8.1.1 A Supplier must take the following actions to enable this outcome:
(c) Ensure awareness and visibility: ensure their staff who have direct contact
with Consumers or former Customers, including personnel working for contractors,
understand the Suppliers Complaint handling process, their responsibilities
under it and are able to identify and record a Complaint.
@@ 8.1.1 A Supplier must take the following actions to enable this outcome:
8.1.1 A Supplier must take the following actions to enable this outcome:
(a) Implement a process: implement, operate and comply with a Complaint handling
process that:
@@
D. requiring Consumers or former Customers to be advised of the Resolution of
their Complaint; and
@@
complaint(X), resolution(X) => [O] informResolution(X)
@@
8.5.1 A Supplier must take the following actions to enable this outcome:
(e) Maintain confidentiality: Suppliers not subject to the requirements of the
Privacy Act must ensure personal information concerning a Complaint is not
disclosed except as required to manage a Complaint with the TIO or with the
express consent of the Consumer.
@@ 8.5.1 A Supplier must take the following actions to enable this outcome:
complaintHandlingProcess(X), personalInformation(X), -subjectPrivacyAct(X) =>
@@
personalInformation(X), requestFromTIO(X) => [O] discloseInformation(X)
@@
The final prompt was evaluated across multiple LLMs. Two diagrams summarize the results: one displaying the standard success scores $S$ (s. Figure 1) and another illustrating the success scores $S^{*}$ under the stricter criterion of perfect formalizations (s. Figure 2). An important observation is that the newly evaluated models generally outperform those considered in [Horner2025]. Specifically, GPT-4.1 yields better results than GPT-4o, and DeepSeek-R1 (0528) surpasses the earlier release of DeepSeek-R1. This trend, however, does not extend to DeepSeek-V3. Moreover, Claude Sonnet 4 demonstrates excellent performance when the Extended Thinking feature is enabled, whereas the same model without this feature yields notably poor results.
@@ 4.2 Consideration of Multiple Law Snippets Simultaneously
In the experiments described in Section 4.1, the prompt was sent together with an individual law snippet to the LLM. This approach ensured optimal focus of the LLM on all details of the individual law snippets but limited the reuse of atom names across snippets. Here, we investigate whether providing multiple law snippets simultaneously enhances formalization performance.
@@ 4.2.1 Incorporating the Formalization History
However, no improvement was observed compared to the single-snippet baseline; in fact, the success scores were marginally lower. A plausible explanation is that the additional context overwhelmed the models, hindering their ability to focus effectively on the current snippet. In particular, the goal of improving the reuse of atoms across multiple law snippets was not achieved.
@@ 4.2.3 Formalizing All Law Snippets in a Single Interaction
Although this approach led to a slight increase in atom reuse across snippets, it exhibited a major drawback: the generated formalizations were often less detailed compared to the baseline obtained in Section 4.1. In particular, important facts were frequently merged into a single rule, even in cases where separate rules would have been necessary for a proper and precise formalization. This issue was especially pronounced for Claude Sonnet 4 with Extended Thinking enabled, though it was also observed in other models.
@@ 4.3 Fine-Tuning
Fine-tuning is a transfer learning technique where pretrained model weights are adapted to a new task through further training. By leveraging knowledge acquired during pretraining, fine-tuning can substantially enhance model performance, particularly in scenarios characterized by limited training data. Prior work has demonstrated the effectiveness of fine-tuning LLMs in improving task-specific outcomes [DBLP:conf/iclr/WeiBZGYLDDL22].
In the present study, fine-tuning experiments were conducted with GPT-4o and GPT-4.1. Due to the proprietary nature of OpenAI’s GPT series, direct access to the model weights is not available. However, OpenAI offers fine-tuning capabilities for selected non-reasoning models through its platform.
Given that only 22 law snippets from the gold standard presented in [DBLP:conf/aicol/DragoniVRG17] correspond to Sections 8.2.1(a)–8.2.1(c), the remaining 44 snippets from unrelated sections were utilized as training data.
@@
Configuration 1 parameters were determined automatically by OpenAI, as recommended for initial fine-tuning attempts [OpenAiFineTuningReference]. However, early signs of overfitting motivated adjustments such as increasing the batch size and reducing the learning rate in Configurations 2 and 3.
After each training epoch, an evaluation has been conducted. The resulting performance is depicted in Figures 5 and 6, where blue bars correspond to non-fine-tuned baselines and other colors represent fine-tuned models. As the first configuration led to suboptimal results, its outcomes have been omitted from the figures for the sake of brevity and clarity.
For GPT-4o, fine-tuning resulted in a slightly improved success score after a single epoch of training under Configurations 2 and 3, relative to the baseline performance of the non-fine-tuned GPT-4o. However, subsequent training epochs led to a decline in performance, indicative of overfitting. Moreover, for GPT-4.1, fine-tuning led to less favorable results compared to using the base model without fine-tuning.
@@
4.4 Two-Stage Pipelines
In this approach, two different two-stage prompting strategies were employed, where the output of the first stage served as part of the input for the second stage. This method aligns with the Layer-Of-Thoughts paradigm, which has been shown to enable complex reasoning in LLMs [DBLP:journals/corr/abs-2410-12153].
4.4.1 Separate Atom Extraction
In this experiment, the intended names of the atoms alongside brief textual descriptions were extracted in the first stage from the legal texts using the system prompt in Listing 3. Three illustrative examples were provided within the prompt, thus applying a three-shot learning strategy. This was followed by the formalization step based on the previously identified atom names.
Extract all the relevant atoms from the legal text in natural language and add a
textual description of them. Each atom should end with "(X)". Do not include
negations in the atom name - these will be introduced later on. Since the law
snippet are talking about complaint handling, in most law snippets, there will be
an atom like complaint(X). Make sure to keep the atoms as simple as possible. If
it is possible to break down the atoms into smaller parts, please do so. For
example, instead of urgentComplaint(X), write complaint(X), urgent(X). The only
exception to this rule is when you can anticipate that an atom will belong into
the consequence. In this case, a longer atom name is better, as each rule can have
only one consequence. Keep in mind that these atoms will serve as antecedents and
consequences in formalized rules - therefore, formalize enough atoms so that
antecedents and consequents can be constructed from them. Formalize at least two
atoms.
@@ 4.4.1 Separate Atom Extraction
8.1.1 A Supplier must take the following actions to enable this outcome:
(a) Implement a process: implement, operate and comply with a Complaint handling
process that:
(v) clearly states that Consumers or former Customers have a right to make a
Complaint and that a proposed Resolution must be accepted by a Consumer or
former Customer before a Supplier is required to implement it;
@@ 8.1.1 A Supplier must take the following actions to enable this outcome:
informRightToMakeComplaint(X): Supplier informs customer of right to make a
complaint.
informComplaintHandlingProcess(X): Supplier informs Customer of Complaint handling
process.
complaintHandlingProcess(X): Supplier has a complaint handling process as per TCPC
section 8.
@@
For each model evaluated in Section 4.1 (except Llama 4 Maverick and GPT-4o mini due to their poor previous performance), we conducted an experiment in this section for the atom extraction step. Moreover, a fine-tuned variant of GPT-4o, specifically trained on examples of atom extraction, was additionally examined. As in Section 4.3, the 44 legal text snippets from [DBLP:conf/aicol/DragoniVRG17] not associated with Sections 8.2.1(a)–8.2.1(c) were used as training data for fine-tuning this model.
In the subsequent stage, the generation of DDL rules was performed based on the legal text and the previously extracted atom definitions. However, we observed that the overall performance remained relatively stable regardless of the model used in the second stage, which justifies limiting the evaluation at this stage of the pipeline to only a few selected models. Based on this observation, we conclude that the primary source of errors lies in the first step of the pipeline. Furthermore, we found that the initial models in the pipeline frequently generate superfluous atoms, which should not be considered by the model in the second step.
@@
Figure 7: Success scores when using a separate atom extraction step
@@
Figure 8: Success scores when using a separate atom extraction step (perfect formalizations only)
The figures indicate that this two-stage pipeline setup performs significantly worse than the new models presented in Section 4.1.
4.4.2 Refinement Step
We used the best formalizations from Section 4.1 as a basis for a refinement step using an additional LLM. The primary objective of this step was to mitigate systematic shortcomings observed in the earlier experiments, in particular the insufficient reuse of atom names across different law snippets. This issue represents a substantial limitation of the prior results (including those discussed in Section 4.2), as consistent atom reuse is essential for obtaining coherent and interpretable formalizations of legal texts.
For illustration, while the gold standard uniformly employs the atom resolvable15Days(X), the LLM-generated formalizations frequently produced a wide range of variants such as resolveIn15Days(X), resolvedBy15Days(X), or cannotResolveIn15Days(X). Such variation introduces unnecessary redundancy and hinders semantic consistency. Moreover, it is undesirable for atom names themselves to contain lexical negations such as “not”, since negation is intended to be represented uniformly by a prefix operator (-) rather than encoded within the atom identifier.
In the refinement step proposed here, all law snippets, together with their corresponding generated rules, were processed jointly by Claude Sonnet 4. The model was instructed to consolidate duplicate or inconsistent atom names into a unified vocabulary, thereby ensuring greater conceptual coherence across the formalizations. The specific prompt used in this stage is provided in Listing 4.
In the following, I will give you an XML file which is composed of mulitple
paragraphs, where each paragraph contains multiple generated rules. Please touch
only the <Generated> element and its children. These generated rules formalize the
law text you find in the <RelevantText> atom using Defeasible Deontic Logic (DDL).
Regarding the syntax: A rule consists of multiple atoms as precondition, which are
seperated by commas, and exactly one atom as conclusion. The conclusion is
separated from the preconditions by a "=>" sign (or =&gt;). Each atom name ends
with (X). Atoms might be negated, which is indicated by a minus sign (-) in front
of the atom name.
However, these formalizations are not perfect. Sometimes, there are (in different
rules/paragraphs) multiple names introduced for the same concept. For example,
"canResolveIn15Days(X)" and "resolvedIn15Days(X)" can be unified. Note that this
is just an example. Please, look out for similar cases where two atoms mean the
same thing and unify them. Pay attention! Look for atoms across ALL paragraphs/
rules that represent the same concept, not just within individual paragraphs. When
unifying atoms that represent the same core action but with different timing or
contextual constraints (for example adviseTimeframeWithin2Days(X) and advise-
TimeframeWithin15Days(X)), you may drop the timing constraint in order to re-use
atoms more effectively. In general, the less atom names you have in total, the
better.
Moreover, if it is not clear from the atom name what the atom is about, please make
the atom name more descriptive. For example, the following atom names are not well
chosen:
* phoneCall(X) - it is not clear what happens in this phone call, better would be
"madeByPhoneCall(X)"
* unresolved(X) - contains a negation, which is not allowed
* acceptance(X) - better would be "customerAccepted(X)"
* complianceClauseC(X) - not clear what clause C is, better would be
"compliedWithClause8_2_1_c(X)"
If necessary, you may also remove atoms or add new ones, but make sure that the
conclusion of the rules contains exactly one atom. Never add new <Rule> elements,
but only modify or delete existing ones.
Make sure that all atom names DO NOT contain negations like "not" or "cannot" -
instead, write a minus sign (-) in front of the atom to show a negation. For
example, if the atom name is "notUrgent(X)", you would write "-urgent(X)" instead.
Note that such negation words can also appear in the middle of an atom name. This
allows you to reuse atoms both in positive as well as in negative contexts. This
should also apply to atom names that you introduce yourself. In other words, in
your output, no <Rule> element should contain an atom name that contains "not" or
"cannot" or something similar. Attention: Do not simply remove the negation word,
but toggle the minus in front of the atom.
When you are done, output the full input I have given to you, but with the
modifications incorporated that you have made. Do not give additional text or
explanations, just the modified XML file.
Listing 4: Prompt for the refinement step
Additional experiments were conducted using alternative models in the refinement step, including OpenAI’s flagship models (o3, GPT-4.1, GPT-4o) as well as DeepSeek-R1. However, these models yielded unsatisfactory results. For instance, some failed to follow the instruction to output the complete XML representation, instead returning only partial fragments of the input. Others generated atom names of excessive length, often extending to 100-150 characters, which rendered the output impractical for use.
The results of this refinement step are presented in Figures 9 and 10.
Figure 9: Success scores when using a refinement step
Figure 10: Success scores when using a refinement step (perfect formalizations only)
4.5 Comparison with [DBLP:conf/aicol/DragoniVRG17]
@@
In this section, we compare our experimental findings with the results presented by [DBLP:conf/aicol/DragoniVRG17]. To ensure methodological consistency, we restrict the comparison to the lower branch evaluation reported in their study. For this purpose, we employ the standard metrics of Precision, Recall, and their harmonic mean, the F1-score, defined as follows:
@@
Our analysis reveals an immediate discrepancy: [DBLP:conf/aicol/DragoniVRG17] report 65 terms and 36 rules in the gold standard, whereas we identified 69 terms and 52 rules across Sections 8.2.1(a)–-8.2.1(c) in the gold standard.
Importantly, in the following analysis, we evaluate the precision and recall of the formalizations produced by the LLMs based on our counts.
@@ 4.5.1 Evaluation of Term Identification
The first level of comparison involves the correct identification of legal terms (referred to as atoms). [DBLP:conf/aicol/DragoniVRG17] reported the following: The gold standard contains 65 terms extracted by the analysts; our system is able to extract 59 terms whose 49 are correct. Therefore, the obtained recall is $90.78\%$ and the precision is $83.05\%$, with a F-Measure of $86.74\%$.” However, it should be noted that these figures are inconsistent. While the precision has been correctly calculated, the recall based on the reported numbers actually amounts to $49/65=75.38\%$ (assuming the stated 65 atoms) rather than the claimed $90.78\%$. Moreover, this discrepancy also leads to a recalculated F1-Measure of $79.03\%$ instead of the reported $86.74\%$. Table 2 summarizes the precision, recall and F1-score values achieved across various configurations and models in our study and compares them with the corrected values reported by [DBLP:conf/aicol/DragoniVRG17]. The highest precision was still achieved by [DBLP:conf/aicol/DragoniVRG17], while the second-best result was obtained by GPT-4o when formalizing all law snippets in a single step. The highest recall was achieved with the additional refinement step described in Section 4.4.2 when GPT-4.1 was employed as the initial model for baseline formalization, whereas the highest F1-score was obtained under the same refinement approach when DeepSeek-R1 (0528) was used in the first stage.
Table 2: Precision, Recall and F1-score for Term Identification
Experiment Model Matched/ Generated Atoms Precision Recall F1
Results from [DBLP:conf/aicol/DragoniVRG17] reported 65 atoms $49/59$ $83.05\%$ $75.38\%$1 0.790.791
counted 69 atoms2 $\mathbf{83.05\%}$ $71.01\%$2 0.770.772
Simple CoI (Section 4.1) GPT-4.1 $59/91$ $64.84\%$ $85.51\%$ $0.74$
Cl. Sonnet 4 (ET) $58/88$ $65.91\%$ $84.06\%$ $0.74$
DeepSeek-R1 (0528) $59/80$ $73.75\%$ $85.51\%$ $0.79$
o3 $58/93$ $62.37\%$ $84.06\%$ $0.72$
o4-mini $58/96$ $60.42\%$ $84.06\%$ $0.70$
GPT-4o $55/77$ $71.43\%$ $79.71\%$ $0.75$
Single Interaction (Section 4.2.3) GPT-4.1 $53/68$ $77.94\%$ $76.81\%$ $0.77$
DeepSeek-R1 (0528) $55/68$ $80.88\%$ $79.71\%$ $0.80$
o4-mini $51/64$ $79.69\%$ $73.91\%$ $0.77$
GPT-4o $53/64$ $82.81\%$ $76.81\%$ $0.80$
Fine-tuned after Epoch 1 (Section 4.3) GPT-4.1 (Config 2) $55/77$ $71.43\%$ $79.71\%$ $0.75$
GPT-4.1 (Config 3) $53/83$ $63.86\%$ $76.81\%$ $0.70$
GPT-4o (Config 2) $56/71$ $78.87\%$ $81.16\%$ $0.80$
GPT-4o (Config 3) $57/72$ $79.17\%$ $82.61\%$ $0.81$
Refinement Step w. Claude Sonnet 4 (Section 4.4.2) GPT-4.1 $60/85$ $70.59\%$ $\mathbf{86.96\%}$ $0.78$
Cl. Sonnet 4 (ET) $58/84$ $69.05\%$ $84.06\%$ $0.76$
DeepSeek-R1 (0528) $58/73$ $79.45\%$ $84.06\%$ $\mathbf{0.82}$
o3 $58/74$ $78.38\%$ $84.06\%$ $0.81$
  • 1
    We recalculated the Recall and F1-score, as the values reported in [DBLP:conf/aicol/DragoniVRG17] are inconsistent. Based on 65 atoms in the gold standard and 49 correct out of 59 generated atoms, the Precision is $75.38\%$ instead of $90.78\%$, resulting in an F1-score of $0.79$ rather than $0.86$.
  • 2
    We counted 69 atoms in the gold standard as opposed to the 65 reported in [DBLP:conf/aicol/DragoniVRG17]. Therefore, we provide calculations based on both figures.